OpenStudy (anonymous):

What is the solution to the following system of equations? y = x2 + 10x + 11 y = x2 + x − 7 A. (−2, −5) B. (2, −5) C. (−2, 5) D. (2, 5)

OpenStudy (solomonzelman):

You are given that \(\color{#000000}{ y = x^2 + 10x + 11 }\) \(\color{#000000}{ y = x^2 + x − 71 }\) and therefore you can imply by a simple substitution that: \(\color{#000000}{ x^2 + 10x + 11= x^2 + x − 71 }\)

OpenStudy (solomonzelman):

Then solve the system by first solving this equation (this will give you the x-coordinate of the intersection), and then plug it into one of the original equations (to obtain the y-coordinate of the intersection).

OpenStudy (solomonzelman):

Note: By definition, intersection(s) ((be it points, lines, or planes in \(\mathbb{R}^n\))) are the solution to the system.